OpenStudy (anonymous):

41k= -2747 ..... help

OpenStudy (anonymous):

Don't let the nagtive scare you., Have to remove the 41 on the left to solve. So we divide both sides by 41: \[\frac{41k}{41}=\frac{-2747}{41}\] \[k=67\]

OpenStudy (anonymous):

divide -2747 by 41 i would use a calculator

OpenStudy (anonymous):

*negative

OpenStudy (anonymous):

k=-2747/41 = -67

OpenStudy (anonymous):

Err -67

OpenStudy (karatechopper):

-67

OpenStudy (anonymous):

Negative signs are my bane. They still kill me in higher math :)

OpenStudy (anonymous):

thanks i had 67 lol not negative but what makes it negative

OpenStudy (karatechopper):

well a neg time a pos is always equalling neg

OpenStudy (anonymous):

ok i get it thanks

OpenStudy (karatechopper):

well a neg time a pos is always equalling neg

OpenStudy (karatechopper):

your welcome